a no. consist of two digit whose sum is 9. if 9 is subtacted from the no.the digit interchange their places.find the no.
Answer by Nate(3500)

Number: 10a + b
a + b = 9 or b = 9 - a
10a + b - 9 = 10b + a
Plug:
10a + b - 9 = 10b + a
10a + 9 - a - 9 = 10(9 - a) + a
10a + 9 - a - 9 = 90 - 10a + a
9a = 90 - 9a
18a = 90
a = 5
b = 9 - a = 9 - 5 = 4
Number: 10a + b = 10*5 + 4 = 54